OpenStudy (anonymous):

solve for y: 2=4x+2y

OpenStudy (anonymous):

2-4x=2y 1-2x=y

OpenStudy (mertsj):

1. Subtract 4x from both sides 2. Divide both sides by 2

OpenStudy (anonymous):

2=4x+2y -4x -4x 2y=-4x+2 2y/2y=4x+2/2y y=-2x+1

OpenStudy (anonymous):

thankss :)